OpenStudy (anonymous):

Find the twentieth term if a36 = 78 and a42 = 90.

OpenStudy (anonymous):

Hey are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you given any other numbers?

OpenStudy (anonymous):

like 1+2+3+..+..

OpenStudy (anonymous):

that is all

OpenStudy (anonymous):

helloo?

OpenStudy (anonymous):

sorry i didnt see your post

OpenStudy (anonymous):

ok so \[a _{n}=a _{1}+\left( n-1 \right)d\]

OpenStudy (anonymous):

\[78=a _{1}+\left( 36-1 \right)d\] \[78=a _{1}+35d\]

OpenStudy (anonymous):

\[a _{1}=35d-78\]

OpenStudy (anonymous):

\[90=a _{1}+\left( 42-1 \right)d\] \[90=a _{1}+41d\] \[90=\left( 35d-78 \right)+41d\]

OpenStudy (anonymous):

solve for d, do you understand?

OpenStudy (anonymous):

168=96d 168/96=d

OpenStudy (anonymous):

humm....

OpenStudy (anonymous):

wow i dont know that is confusing

OpenStudy (anonymous):

\[a _{n}=a _{1}+(n-1)d\] \[a _{20}=a _{1}+(20-1)d\] \[a _{20}=a _{1}+19d\] \[d=168/96\]

OpenStudy (anonymous):

is that final answer

OpenStudy (anonymous):

are you sure thats all your problem says?

OpenStudy (anonymous):

yes